OpenStudy (anonymous):

find the value of a: 5rt5 * 5^3/ 5^(-3/2) = 5^(a+2)

OpenStudy (anonymous):

\[5\sqrt{5}*5^3/3^{-3/2} = 5 ^ {a+2}\] \[a + 2 = 3/2+3/2 + 3 = 6\]\[a =4\]

OpenStudy (anonymous):

can u explain it in steps please

OpenStudy (anonymous):

sorry dainel u had written wrong question but the answer is correct'

OpenStudy (anonymous):

I don't think I wrote the wrong question... did I?

OpenStudy (anonymous):

thanks for ur answer /3−3/2

OpenStudy (anonymous):

there should be / 5^(-3/2)

OpenStudy (anonymous):

If a and b are positive integers such that ab = 125, then (a-b)^(a+b-4) is equal to ?